term HOWTO

Bill Reynolds bill@goshawk.lanl.gov, Patrick Reijnen patrickr@cs.kun.nl September 3, 1994

Contents

1 Introduction 2 1.1 About This Document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 What is term? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 How term works 3 2.1 Nomenclature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Setting Things Up 4

3.1 Build term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.1 Build term upto version 1.15 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.2 Build term, versions 1.16 upto 1.19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.3 Build term, versions 2.0.0 and higher . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.2 Setting environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Test term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.4 Make a Transparent Link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.5 Run linecheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.6 Try Running term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.7 Terminate your term connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.8 Gotchas to Watch Out For . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 term clients 12

5 X and term 12

6 tredir  13

6.1 tredir can bite! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 6.2 Stupid tredir tricks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 6.2.1 X windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

6.2.2 Mail with term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

7 Automating Things 19

8 Other programs and term 21

8.1 Use ready made binaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

9 Term speed table 21

10 Term clients used in the world 22

11 Things to remember 23

12 Other Stuff 24

13 Acks 24

1 Introduction

1.1 About This Document

This HOWTO attempts to clear up some of the confusion of using term, Michael O’Reilly’s remarkable program that allows you to multiplex your serial line and set up an ersatz network connection. By and large, the documents that come with term are quite good, and this HOWTO is not intended to replace them. The intent of this document is to give some background on how term works and detail the steps in getting some of the more common networking services working under term. I should emphasize that this document does not cover everything there is to know about term. After reading it, you should read the term manpages, since they include information not contained here.

1.2 What is term?

term is a program, written by Michael O’Reilly (michael@iinet.com.au), that is run over a serial line to allow multiple connections to operate concurrently - i.e. you can be downloading a file over your modem while working on a (different) remote system via the same modem connection. term can also be used to open up X client windows over a serial connection. Through the tredir utility and the tudpredir utility in term versions 2.0.x, term can provide many of the “traditional” network services: mail, news, ftp, telnet, xarchie, etc. In a sense, term is very much like other serial protocols such as SLIP or PPP. term’s advantage is that it can be run entirely from user space, requiring neither kernel support nor support from system or network administrators.

2 How term works

2.1 Nomenclature

I am assuming you are dialing up a system through some sort of terminal server. I use the terms “local” and “remote” to refer to the home and network connected systems respectively (unless I use them to mean something else :-).

term provides the local machine, which has no network connection, but is connected via a serial line to a remote machine which is in turn connected to a network, with network services. Let’s look at how a machine with a “traditional” network connection provides these services. First the user invokes a program, like ftp or telnet, that requests a network service. What these program do is make a system call requesting network services. The operating system then obtains these services via its network interface (e.g. it sends and receives packets over the ethernet). SLIP and PPP do exactly this, by converting your modem line into a network interface, which is in principle no different from an ethernet. The downside of this is that these protocols make the modem-connected machine part of the network, just like any other machine. This entails all the administrative burdens associated with being a network node (more actually, since the modem link must also be administered).

In the absence of a network connection like SLIP or PPP, what does one typically do? Well, you dial up your network connected machine, read your mail, your news etc, if you need a file, you first transfer it to the remote machine and then download it to your local machine using kermit or some other communication program. This is a bit of a pain, especially since you can only really do one thing at a time that uses your modem link. The idea behind term is basically to automate and multiplex this process. term is invoked on both the local and remote machines, and the two processes communicate with one another over the modem line. When you need a network service, you make a request to the local term daemon, which forwards the request to the term daemon on the remote, network-connected, machine. The result is then returned over the modem line.

To be more concrete, say you want to retrieve a file by ftp. First you need a version of ftp that can speak to term. You invoke this termftp as you do a regular ftp, say termftp nethost.gov, but this special version makes its network request to the local term daemon instead of the kernel. The local term forwards this request, over the modem line, to the remote term, which opens an ftp connection to nethost.gov, and transmits the data back over the modem link. term is smart enough to have many different things going on at once, so you can have several different network sessions using the same modem link, e.g. you can be logged into another distant host via termtelnet while the termftp transfer is going on.

If this is too abstract (or unclear) do not worry; the important piece of information to get out of this section is that there are two copies of term running, one on each end of the modem link.

3 Setting Things Up

3.1 Build term

When you are lucky, this should just involve a make. Most probably however, you need to do more. Due to new features in newer versions of term (sharing, configure) it is now a bit more complicated to create obtain your term binary. A couple of ways can now be followed to obtain your binary.

To cover all these ways term can be build I will split this section into three parts:

1. Build term upto version 1.15

2. Build term, versions 1.16 upto 1.19

3. Build term, versions 2.0.0 and higher

3.1.1 Build term upto version 1.15

For these versions of term building should invoke no more than the commands

make DO=install OS-type
make installman

You will find term, its clients and the manual pages nicely built and installed and ready for use after this.

The only thing you may want to do is changing some of the pathes in the Makefile or changing some of the compiler flags.

3.1.2 Build term, versions 1.16 upto 1.19

Starting with version 1.16 the concept of sharing the term connection with other users has entered term. With this concept built in also the ways of building term are extended.

To build term you can now choose one of the following ways:

1. You are an ordinary user (no root acces) and you do not want to share the term connection with other users. As user who does not want to share the term connection with other users you should do the following to build term:

     make DO=install OS-type
     make installman

After this term, its clients and the manual pages are built and installed.

2. You are an ordinary user (no root acces) and you want to share the term connection with other users. As user who wants to share the term connection you should do the following:

make DO=installshare USERSHARE=$HOME/term OS-type
make installman

After this term, its clients and the manual pages are built and installed. Furthermore, you will have a directory $HOME/term with permissions drwxrwxr-x

3. You are root and you do not want to share the term connection with other users. As root who does not want the term connection be shareable you should do the following to build

term:
make DO=install OS-type
make installman

After this term, its clients and the manual pages are built and installed.

4. You are root and want to make the term connection shareable.

Create a term group with no member by adding the following line in /etc/group:

term::16:root

or any other unused GID than 16 when 16 is already in use. Now do

make DO=installshare OS-type
make installman

After this term, its clients and the manual pages are built and installed. Furthermore, you now have a directory called /usr/local/lib/term (default) owned by group term and with permissions drwxrwxr-x

3.1.3 Build term, versions 2.0.0 and higher

For term versions 2.0.0 and higher there are many ways to build the term binary and the clients. All of these can be done both by root and by ordinary user:

1. Build term in private mode without full term networking

2. Build term in private mode with full term networking

3. Build term in shared mode without full term networking

4. Build term in shared mode with full term networking

In these versions of term also the configure concept is used. When configure is run it checks on what operating system you are trying to install term, whether the source directory is available or not and if any runtime options are set. According to the things found configure then creates a Makefile using Makefile.in which is provided in the term package.

Two of the more important options to configure are --root and --user which state whether term will be installed by root or an ordinary user.

Build term in private mode without full term networking To build term this way you need to execute the following commands (both for root and ordinary user):

./configure --root  OR --user
make install installman

This builds the binaries and installs these binaries and the manual pages.

Build term in private mode with full term networking. To build term this way you need to execute the following commands (both for root and ordinary user):

./configure --root  OR --user
make installnet installman

This builds the binaries and installs these binaries and the manual pages.

Build term in shared mode without full term networking To build term this way you need to execute the following commands (both for root and ordinary user):

./configure --root  OR --user
make share installman

This builds the binaries and installs these binaries and the manual pages.

Build term in shared mode with full term networking To build term this way you need to execute the following commands (both for root and ordinary user):

./configure --root  OR --user
make share installnet installman

This builds the binaries and installs these binaries and the manual pages.

3.2 Setting environment variables

term knows a couple of environment variables which can be set by users. The first three of these I will explain are

• TERMDIR

• TERMSHARE

• TERMMODE By setting these variables you can control the way term is run.

For term versions upto 1.15 only the variable TERMDIR is important (these versions do not know the shared mode). For these versions TERMDIR should be set as follows:

setenv TERMDIR $HOME     (csh or tcsh)
export TERMDIR=$HOME     (bash)

Starting with version 1.16 term also lnows the variables TERMSHARE and TERMMODE. With these variables term can be told to run in private mode or in shared mode. I will explain how to set the variables for private mode and shared mode.

1. Running term in private mode can be done setting the variables TERMDIR and TERMMODE in the following way: For csh or tcsh

setenv TERMDIR $HOME
setenv TERMMODE 0

For bash

export TERMDIR=$HOME
export TERMMODE=0

2. When you want to use term in shared mode there are two ways of setting the variables:

(a) When term is installed as a SUID program only TERMMODE has to be set (?? README in the term package).

setenv TERMMODE 'number'    (csh or tcsh)
export TERMMODE='number'    (bash)

For 'number' you have to enter 1 when you are using one of the term versions 1.16 upto 1.19 and 2 when you are using term version 2.0.0 or higher.

(b) When term is installed as a SGID program the variables have to be set in the following way: For csh or tcsh

setenv TERMMODE 1
setenv TERMDIR /usr/local/lib/term
setenv TERMSHARE $TERMDIR

For bash

export TERMMODE=1
export TERMDIR=/usr/local/lib/term
export TERMSHARE=$TERMDIR

Setting the variables this way makes it possible to start old clients (clients linked to an older version of client.a) in shared mode.

Starting with version 2.0.0 term also knows the variable TERMSERVER. You need to set this variable when you have multiple modems and you have more than one conection at a time. To specify which connection to use, you must start term with a server name:

nohup term -v /dev/modem1 Connection1 &
nohup term -v /dev/modem2 Connection2 &

Users should then set the variable TERMSERVER to the connection name they want to use:

setenv TERMSERVER Connection1   (csh or tcsh)
export TERMSERVER=Connection2   (bash)

3.3 Test term

Do a make test to build term’s test daemon. Test works by running two copies of term on your system, a “local” and a “remote” copy. Both of these will read your termrc, so you can adjust their behaviour. Run test as ./test so as to avoid your system’s test. You should now be able to do a trsh and a tupload (try tupload ./term /usr/tmp - you should get a copy of the term binary in /usr/tmp ). The local term’s output should show up in local.log, the remote’s in remote.log. You can start term up with a -d255 flag to enable debugging output to be written to these files, or enable debugging in your termrc.

3.4 Make a Transparent Link

Presumably, you can establish a modem connection between your local and remote hosts. Typically, your dialing into some kind of terminal server and connecting to your remote host from there. You are also using some kind of terminal software, such as kermit or seyon to talk to your modem (the examples in this document will use kermit, since that is what its author uses). If you are having trouble with your modem, or your terminal software, take a look at the Serial-HOWTO, that should help you out.

Having established your link, you want to make it as transparent as possible. Check the commands on the terminal server (help or ? will usually get you started). Go for the 8 bit options whenever possible. This may mean changing the way you log in to a system, e.g. if the server uses rlogin, you may have to use it and give it the -8 flag to make it transparent. Especially watch out for xon/xoff flow control, you do not want that, try to enable rts/cts (hardware) flow control. You may need to check your modem docs to learn how to configure it to do 8-bit rts/cts communications.

3.5 Run linecheck

linecheck is program that is supplied with term. It checks the transparency of a link, providing configuration information that term needs to run correctly. linecheck will send each of the 256 possible eight bit characters over the link and verify that each was transmitted successfully. term needs to be configured to deal with characters that cannot be transmitted over the link, and linecheck determines what characters these are. You use linecheck after you’ve established as transparent a modem link as possible. To run linecheck, do the following

  1. On the remote system, run linecheck linecheck.log.

2. Escape back to your local system and suspend your comm program (^Z under kermit) (otherwise it will steal characters from linecheck).

  3. On the local system run linecheck linecheck.log > /dev/modem < /dev/modem.

When linecheck is done, you will find a set of numbers at the bottom of the linecheck.log files. These should be escaped in the termrc at the other end of the link. For example, on my system, my local linecheck.log said nothing and my remote linecheck.log said to escape 29 and 157. Therefore, my local termrc escapes these characters and my remote termrc escapes nothing. If I escape a character on one end, I have to also ignore it on the other, so, in this example, I shall have to ignore 29 and 157 on my remote system.

If linecheck hangs, try using

linecheck linecheck.log 17 19 > /dev/modem < /dev/modem

this will escape your xon/xoff (flow control) characters, which will hang your line if you have got software flow control. If this solves the hanging problems, you will want to escape /ignore 17/19 in both termrc’s. If your terminal server has other demonic characters that will shut it down, try running linecheck with those characters escaped, as above. You can spot these characters if linecheck hangs. If this is the case, kill it, then look in the logfiles. The last characters transmitted are likely to be the culprits. Try it again with these characters escaped.

In summary, my local termrc has the lines

escape 29
escape 157

and my remote termrc has the lines

ignore 29
ignore 157

since my remote linecheck.log said to escape 29 and 157.

3.6 Try Running term

Log into the remote system, making the link as transparent as possible (if you have not already). Fire up term on the remote end. I use the following:

exec term -l $HOME/tlog -s 38400 -c off -w 10 -t 150.

Let’s run down each option one by one (note that I could just as easily have put these options in my termrc. I did it this way because it saves editing a file while getting term set up). exec means to destroy your current shell, running the given program in its place. I exec things because I do not intend to use my login shell again, so it is just wasting memory. If you are debugging the link and can reliably kill the remote term, you might not want to do an exec.

-l $HOME/tlog. This logs errors to the file tlog in my home directory. Very useful for debugging. No reason not to do this.

-s 38400 : I have got a 14400 baud modem, with compression. For optimal compression ratios, I want to be able to push bits down the pipe as fast as possible. For a slower modem, you should use something lower. Note that if you have a slower machine with 16450 uart on your serial port, high baudrates can cause data loss by overloading the chip on your serial port. term will recover from this, but if you see a lot of error messages in your log file, (or get overrun warnings from linux kernel versions 0.99pl15 and up) you again might want to lower this number.

-c off : This turns data compression off. I have got a compressing modem, and I do not want to compress things twice.

-w 10 -t 150 : Again, these are options to optimize my fast modem link. I have set my window to 10 and my timeout to 150. This is per the recommendation in the term setup man page.

Escape back to your local machine and suspend your comm program (^Z under kermit). You do not want it running while term is running, because it will fight with term over the serial port. If you can convince your modem to not hang up when you exit your comm program (when it toggles DTR), you could just exit the program at this point.

Now run term locally. I use:

term -r -c off -l $HOME/tlog -s 38400 -w 10 -t 150 < /dev/modem > /dev/modem &

The only option here that is different from the other end is the -r option. You need this on one end, otherwise term clients will spontaneously die on you. I need to tell term where the modem is, so I point both standard input and output at /dev/modem (that is what the < and > do). I also run it in the background, so I can use this screen for something else if I want to.

term should work now :-). Try a trsh, and see what happens. If you hang, or your link seems slow, take a look at your tlog on each end. Are you getting timeouts or error messages? If so, then you have configured something incorrectly. Try again (after you have finished reading this :-). Note that the connection won’t seem blazingly fast, especially if you are using compression - it will be a little jumpy. The real speed comes in during file transfer and the like.

3.7 Terminate your term connection

Most certain, after you have done a lot of work using term, you want to finish your work and bring your term connection down. For this to be realized there are three ways:

1. Kill the term programs on both sides of the connection. This is the least recommended way of terminating your connection

2. A better way is to execute the following command locally:

     echo '00000' > /dev/modem

This wil nicely terminate your term connection. It will work for all version of term. Make sure that the sequence of zeros contains at least five zeros.

3. Starting with version 1.14 there is the program tshutdown (actually for version 1.14 it is available as a patch, for newer versions it is in the package). Executing tshutdown nicely terminates your term connection.

3.8 Gotchas to Watch Out For

If your .term or term directory is mounted with NFS you need to set the flag -DTERM NFS DIR in the CFLAGS line of the Makefile. Alas, for the author this flag causes a compile error when term 1.19 is compiled on a machine running sunOS 4.*.

Are you running the right binary? term has been updated quite a lot, and many systems have different versions of the programs floating around. Make sure you are using the right version. Note that this applies to linecheck too. You can use bash’s type -a or the whereis command to find which program is being run. term versions after 1.11 should print out their version number when they start up. (Although version 1.14 claims to be 1.12. Sigh.)

Do you have the right termrc? Depending of the version of term you are running and the way you installed term (being root or user) this file has to be in one of the following directories: /usr/local/lib/term, $HOME/.term or $HOME. Some systems have pre-installed termrc’s, make sure they are gone before you set things up. If you are running things as root, lookout for /.term. term creates files (sockets actually) while it is running, so it has its own directory, $HOME/.term, where the file termrc goes (note, there is no leading dot in termrc!).

With newer versions of term the structure of the directory tree under /usr/local/lib/term has changed a couple of times. If you are not aware of this, it can cause all kinds of error messages. The best thing to do is to delete the directory tree under /usr/local/lib/term (save your termrc) and then install your new term version. This way, you avoid the struggle with a messed up directory tree.

Once you have got term running, you might want to try to get things optimized. A good way to measure the speed of your link is to run tmon in one window while up/downloading a file in another. Try both (big) text files and compressed files, the plain text should go a factor of two-ish faster than the compressed files. The parameters you want to fiddle with are compression (-c), windows (-w), timeout (-t) and retrain (-A. Watch out with the retrain parameter. I got a performance decrease of 80% to 90% compared to running term without the retrain parameter. It is not clear if this is a bug in term version 1.19, the version I am running and if this problem exists only with term version 1.19.

Compression: you want this on if you do not have a compressing modem. If you do have such a modem, turn compression off, otherwise you will be compressing things twice, which typically increases the amount of data transmitted. Compressing modems are ones that use the MNP-5 or V42.bis protocols. Check your modem documentation and the message when your modem connects.

Windows: this is the number of chunks of data, or packets, that term will let go over the line before it gets an acknowledgment (or ack) from the remote term. For fast modems, increasing this can be a win, for slower links this can overwhelm the remote end.

Timeout: the time term will wait for an ack. If you have increased windows, and you are getting timeouts in your logfile, try increasing this.

For 14400/V42.bis, I use -c off -w 10 -t 150. I typically get around 1700 cps on compressed files and 3500 cps on ascii files using tupload.

4 term clients

term provides several default clients. They include trsh, tmon, tupload, tredir, txconn and in newer versions trdate, trdated and from version 2.0.0 tudpredir. This section will deal with trsh, tmon, tupload, trdate and trdated. The others each have their own section. No term client will work until you have established a term link.

tmon is a simple utility to monitor the statistics of your link. It prints a time histogram of characters transmitted and received. It is invoked simply as tmon. Since around version 1.11, tmon has had a bug that causes some information to be garbled. Send me mail if you’d like a patch to fix this.

trsh is similar to rsh. Without arguments, it spawns an interactive shell on the remote system (i.e. it logs you in). trsh is one of the primary means of accessing the remote end of the link via term. If given an argument, trsh executes that argument as a command on the remote system. For example trsh ls would give you a listing of the files in your home directory on the remote system.

tupload is the term’s basic file transfer program. It will transfer a file, given as its first argument, to the opposite end of the link. By default, the files will be put in the same directory that you invoked term from on the other side. To put files in another directory, give its name as a second argument to tupload. For example, if I want to put a copy of the file term114.tar.gz in /usr/tmp on the remote system, I would type tupload term114.tar.gz /usr/tmp. There is no client to do a download - to download files, trsh to the remote system and do a tupload.

trdate is a time setting utility. it reads the time on the remote machine and sets the local clock to the remote time. It must be run as root.

trdated is the daemon version of trdate. When it is started in rc.local it is run as daemon in which case it updates the time every 5 minutes. Even when there is no term connection, this daemon will start up when set in the rc.local. Once a term connection is created it starts updating the time.

5 X and term

term allows users to open up X windows on the local machine from clients that are running on a machine on the network. This is done by using the txconn client. txconn is executed on the remote, network-connected machine, it is invoked simply as txconn. It goes into the background and returns a number on the standard output, this number is the display number that clients should use to access the X server on the local machine. An example should make this clear. I am logged in, via trsh, to my remote term host, named foo. On foo, I do the following

6. tredir  13

foo$ txconn
Xconn bound to screen 10
:10
foo$

Now, on any host that I wish to run an X client, that is to display on my local machine’s X server, I do setenv DISPLAY foo:10 (for bash you should use export DISPLAY=foo:10). In some cases it can furthermore be necessary to do a xhost + foo on your local machine. Now when I start the client, it will try to connect to screen 10 on machine foo, but txconn is listening to this screen, and will forward all X protocol packets via term to the X server on the local host - i.e. the window will open up on your local machine.

It is possible to go the other way - run a client on your local machine and have it open up a window on a remote machine on the network, however we will defer explaining this until after we have discussed tredir.

txconn is not terribly secure, anyone can connect to your local server via term and perform all sorts of mischief. If you are worried about this sort of thing, it might be a good idea to consider using xauth to authorize your connections. Consult the xauth man page.

The X protocol is not very efficient, it wastes some bandwidth. This is usually not a problem over an ethernet, but can be murder over a modem. X11R6 is supposed to introduce a low bandwidth version of the X protocol, LBX. However, this is not of much use, since R6 has not been released at the time of this writing. There is a utility named sxpc which compresses the X protocol, improving response over serial lines. sxpc includes a write-up on how to get it working with term, and is recommended. The sxpc package also explains how to use xauth, so it is doubly recommended.

6 tredir

tredir is one of term’s most powerful utilities, allowing most important network services to be performed over a term link. Before we explain how to use tredir, it is necessary to give some background on network services. We have talked about network services before, but we have not said exactly what they are. Services are just that - services that are provided by the network. Examples of services include telnet, which provides logins between machines, the File Transfer Protocol, ftp, which transfers files between machines, and smtp, the Simple Mail Transfer Protocol, which is used whenever you send electronic mail. Each network service has a port number associated with it. The mapping of port numbers to services is given in the file /etc/services. This file should be the same on all internet-connected machines.

How are these services invoked? Each networked machine runs a daemon called inetd, which listens for attempts to connect to the network ports. These requests can come from either the network or the local machine. A network service is obtained by connecting to a particular inetd port. When a network request is made, inetd knows exactly which service is involved by the port number that request is made on. If inetd is configured to do so, it then provides the relevant service to the requesting connection. inetd’s configuration is given by the file /etc/inetd.conf, which has a list of the services that inetd provides. For more information, see the man pages for inetd and inetd.conf.

You can communicate directly with network services by using telnet (n.b. not termtelnet). For example, to talk to the sendmail, (or smtp) daemon on machine machine name, you can do a telnet machine name smtp, or telnet machine name 25, (since 25 is the number assigned to smtp in /etc/services). You should get a polite greeting from the daemon on the remote machine. This is a very useful trick for debugging network problems and checking ports redirected with tredir (see below).

tredir works very much like inetd. It runs in the background as a daemon, listening to the network ports, waiting for a request. When a request for a service is made, instead of providing that service, as inetd does, tredir forwards the request over the term link to the remote term, which makes the request over the network, returning the result back over the link to the local client. tredir can forward the request to any machine on the network, but by default sends it to the machine on the other end of the term link. tredir “redirects” network services over the term link (hence the name “tredir”).

An example should make this clear. Let’s redirect a local port to the telnet port on the remote machine. To do this we would do tredir 2023 23. Now, anyone who connects to port 2023 on the local machine will be redirected to port 23 (telnet) on the remote machine. Here’s an example session, the local machine

is mymachine.modem.home and the remote machine is netsun.

$ tredir 2023 23
Redirecting 2023 to 23
$ telnet localhost 2023
Trying 127.0.0.1...
Connected to mymachine.modem.home
Escape character is '^]'.

SunOS UNIX (netsun)
login:

This example is actually quite useful. If I were instead to do the tredir on netsun, I could then telnet in to mymachine from the network, simply by connecting to the redirected port on the networked machine (using

telnet) - i.e. telnet netsun 2023.

The general principle in using tredir is to redirect the desired service to a machine on the network. Our next example will allow us to read news on the local machine over our term link from a news server on the network. News is provided by the nntp service, port number 119. All decent news readers allow you to specify what port number they will use, either via a configuration file or an environment variable. Let’s specify this local port to be 2119. Now, let’s say that our news server is news.domain.org. We will redirect port 2119 to port 119 on news.domain.org, we will then tell our news reading software that the nntp server is located at port 2119 on the local host (??). Since this will depend on the news reader that you use, I will just test the link with telnet instead of firing up a news reader:

$ tredir 2119 news.domain.org:119
Redirecting 2119 to news.domain.org:119
$ telnet localhost 2119
Trying 127.0.0.1...

Connected to mymachine.modem.home.
Escape character is '^]'.
200 news.domain.org InterNetNews NNRP server INN 1.4 07-Dec-41 ready
(posting ok).

If you can get this far, all you have to do is configure your news reader to be able to read news via term. (n.b., if you read news like this, be sure that in all your posts you set a Reply-To: header to an network email address that you can be reached at, otherwise people who want to get in touch with you will be sending mail to whatever (wrong) data your newsreader puts in the From: header).

6.1 tredir can bite!

The astute reader, after reading the last example will be wondering why port 2119 was redirected to port 119 - since newsreaders default to port 119, why could not I do a tredir 119 news.domain.org:119 and skip the newsreader configuration? The answer is that all ports numbered less than 1024 are “reserved ports”, and only the superuser can listen on them. If one is willing to take a security risk and make tredir an suid program, or run tredir as root, then one can redirect reserve ports and avoid the hassle of renaming services.

Another problem with using reserved ports is that inetd is often already listening to these ports, and only one program at time can listen to a port. In order to use such a port, you must change inetd.conf so that inetd no longer listens on the port you want to redirect. This is most easily done by commenting out the line with the offending service by putting a # character at the beginning of the line. The superuser must then send inetd a HUP signal (kill -1 inetd-pid) to get it to reread its configuration.

All of the services we have been discussing so far are TCP (Transmission Control Protocol) services. TCP is just a protocol for sending packets (data) over a net. term versions upto 1.19 only support TCP. Starting with version 2.0.0 term also supports UDP (User Datagram Protocol). This makes things like NFS and archie available over a term link.

6.2 Stupid tredir tricks

In this section we will describe some of the more common uses for tredir. We have already described how to redirect nntp and telnet services, here we will give some more complicated examples.

6.2.1 X windows

In a previous section, we described how to get an X client running on the network to open a window on your home machine using txconn. The same technique could be used on your home machine to display a client on the machine at the remote end of your term link. But how does one display an X client on a network machine that is not the remote end? The answer lies in knowing that X uses a particular network service just like the other programs we have been discussing. An X server listens for network request on a port whose number is given by the formula port = 6000 + displaynumber, e.g. an X server managing screen 0 on a machine would listen to port 6000, if it were managing screen 2, it would listen on port 6002. When you set your DISPLAY environment variable to xmachine:n, your X clients will try to connect to port 6000 + n

on xmachine.

We can use this to trick X clients on your local machine to open up windows on remote displays. Let’s say I want to open up an xterm, running on my local machine, on display 0 of machine xmachine, which is running someplace on the network. I first pick a local display number, say 2 (do not use 0, since that is what your local X server will be using). I will map this display to display 0 on xmachine. In terms of ports, this means I want to redirect the local port 6002 to the remote port 6000. I do the following

$ tredir 6002 xmachine:6000
$ setenv DISPLAY localhost:2
$ xterm

This should open up an xterm on machine xmachine. Note that I set the DISPLAY to localhost:2. This is because X clients will sometimes use unix domain sockets instead of internet domain sockets, at their own option, when connecting to a local display, if DISPLAY is set to :2. localhost:2 says to use a tcp connection.

Note that as far as xmachine is concerned, the X request is coming from the machine on the remote end of your term link (remotemachine) - so if you need to authorize the connection, you should either do an

xhost + remotemachine on xmachine or use xauth to update the .Xauthority file on your local machine

for display number 2, using the key from xmachine.

Again, to speed up X connections, you can use the program sxpc, which includes an explanation of how to use tredir to establish the link and authorize it using xauth.

6.2.2 Mail with term

Well, you asked for it. Electronic mail has the justifiable reputation of being one of the most difficult things to get working right on a UNIX system. To really get term working correctly with mail means that you have to understand how mail works, which is beyond the scope of this document. To learn more about mail, you should consult a book on UNIX system administration and/or the comp.mail.misc FAQ, available for anonymous ftp on rtfm.mit.edu in pub/usenet/comp.mail.misc. There are also currently two packages available for anonymous ftp on sunsite.unc.edu that will help you get mail running under term - they are

term.mailerd+smail by Byron A. Jeff and the BCRMailHandlerXXX by Bill C. Riemers.

That being said, we will give a thumbnail description of how mail works. There are two parts to getting mail running, sending messages and receiving messages. We will begin with sending messages from your local box to the network.

There are two classes of mail programs. The first is the mail user agent (MUA). MUA’s help you read, compose and send messages. Examples of MUA’s are elm, pine, Mail and vm. MUA’s do not really do any networking, they just put the messages together - the real work of sending mail is done by the second class of mail programs, the mail transfer agents (MTA’s). These are invoked by the MUA’s. They take the message, decide where to send it by looking at the address, and then actually deliver it over the network.

The two most common MTA’s on Linux systems are sendmail and smail. The basic idea is to get your MTA to connect to another MTA running on a machine on the net that will know what to do with your message. This is done by redirecting a local port to the smtp port on the net machine. You then have to tell you MTA to take any message it does not know what to do with and send it out over the redirected port on your local machine to the MTA on the remote machine, which will then route your message to its correct destination.

How do we do this using smail? We first redirect a port to the smtp port on the network mail machine

(mailhost):
tredir XXXX mailhost:25

here XXXX is the port number that the smail on the localhost will connect to (note that I have to give this port a name in my /etc/services to get smail to recognize it). smail has several configuration files that

usually live in /usr/local/lib/smail. The ones we care about are config, routers and transports.

Note that I am assuming you have already got smail configured correctly for local mail - delivery to files and pipes and such. Again, consult the documentation if you do not.

In the file config, we put the following definition:

smart_path=localhost

localhost is the machine that smail connects to when it does not know what to do with a message.

In routers we put

smart_host:
driver=smarthost,
transport=termsmtp;
path = localhost
In transports we put

termsmtp:        driver=tcpsmtp,
        inet,
        return_path,
        remove_header="From",
        append_header="From: YOUR_NET_ADDRESS",
        -received,
        -max_addrs, -max_chars;
        service=YOUR_SMTP_SERVICE,

In the above, the header lines change the From header in all your outgoing mail to the address, YOUR NET ADDRESS, which is the network address you want mail for sent. If more than one user is going to be using your term link, you will have to do something more fancy, like keep a database of local user’s network addresses and insert these in the From: headers.

The service line is the name of the local portnumber that you have redirected to the smtp port on the network connected machine. On my version of smail, I cannot just set this to a number, I have to set it to a name, like “foo”, and then define “foo” in my /etc/services to be the number of my redirected port. If you use an suid tredir and just redirect the smtp port (25), you do not need to define this.

This should be enough to get you going. If you decide to use sendmail, the principles are the same but the details differ. Ronald Florence (ron@mlfarm.com) told me that the stock Sun sendmail will not send multiple queued messages over a redirected port; BSD sendmail 8.6.9 works fine. He made the following changes to sendmail.cf to get it working with term. In his case, the default sendmail port (25) is used for SMTP traffic over a local ethernet, so Internet mail is forwarded to a redirected TCP port.

#
#Create the termsmtp mailer, which sends mail via a re-directed TCP port
#
Mtermsmtp,P=[TCP], F=mDFMuCXe, S=22, R=22, A=TCP $h PORTNUMBER

Here, PORTNUMBER is the number of the redirected port on the local machine. This should be an unused port over 2000. We next tell sendmail to which machine to connect to, and set termsmtp as the default mailer.

#
# major relay mailer
#
DMtermsmtp
#
# major relay host: use the $M mailer to send mail to other domains
#
DR HOSTNAME
CR HOSTNAME

Here HOSTNAME is the name of your local host (does localhost work?). The last entry goes under Rule 0 to forward Internet mail.

# Pass other valid names up the ladder to our forwarder
R$*<@$*.$+>$*           $#$M    $@$R $:$1<@$2.$3>$4     user@any.domain

When the term connection is established to the Internet host, run the following commands on the local machine.

tredir PORTNUMBER internet.host:25
/usr/lib/sendmail -q

We now turn to receiving electronic mail using term. We will assume that mail is sent to your account on the network machine mailhost. The simplest solution is to just use trsh or termtelnet to log on to mailhost and read your mail on there. However, it is also possible to have your mail automatically downloaded to your local machine. One way to do this is to use the Post Office Protocol, (POP). POP was designed for exactly this purpose: to deliver mail to machines that have intermittent network connections. To use POP, you must have a POP server installed on mailhost. Assuming that you do, you can then use a POP client to download your mail every few minutes. This is done, as you might expect, using tredir. The POP service is 110 (note that there is an older protocol, POP-2, which uses port 109; in this document we describe POP-3, which is the latest version of POP). There are several POP clients available. One, written using the script language perl is pop-perl-1.X, written by William Perry and maintained by myself - it can be found on

sunsite in /pub/Linux/system/Mail.

To use POP, you redirect a local port to port 110 on mailhost and configure your client to retrieve your mail from localhost using the local port. As an example, we will assume that there is a POP server running on mailhost. We will redirect the local port 2110, and fire up the pop-perl client:

$ tredir 2110 mailhost:110
Redirecting 2110 to mailhost:110
$ pop
Username: bill
Password: <enter your password for mailhost>
Pop Host: localhost
Pop Port: 2110
Starting popmail daemon for bill

If you do not have a POP server available, the BCRMailHandler package has a program to download your mail over a term link to your local machine. I have not used it, but anyone who has is welcome to comment.

7 Automating Things

Now that you know how to get all of your network services over term, it would be nice to set things up so your link is set up and configured automatically. There are basically an infinite number of ways of doing this, depending on what comm program you use and how you log in to your remote system.

One program, that I have not used, but I have heard is quite nice, is fet: a front end for term. It is designed to log you into a remote system and fire up term and all your tredir’s. Any comments on fet would be most welcome.

I shall now give an example of a set of commands that use kermit to log into the remote system and then performs all of the term initializations. Obviously, if you use these examples, you will have to modify them for your own login procedures.

The command which is actually invoked is the shell script knet, given by:

#!/bin/sh
/usr/bin/kermit -y $HOME/.kerm_term > $HOME/klog < /dev/null 2>& 1
exec $HOME/bin/tstart >> $HOME/klog 2>& 1

The script .kerm term is given by:

pause 2
# The number you want to dial
output atdtXXXXXXX \13
# Login to the terminal server
input 145 {name: }
output MYNAME \13
input 3 {word: }
output MYPASSWORD \13
input 5 {xyplex>}
# Make the line transparent
output term telnet-t \13
output term stopb 1 \13
# Connect to the remote host
output telnet remotehost.somedomain.org \13
input 10 {ogin: }
output MYOTHERNAME \13
input 3 word:
output MYOTHERPASSWORD \13
pause 5
# Fire up term on the remote host
output exec term -s 38400 -l $HOME/tlog -w 10 -t 150  \13
! /usr/bin/term -r -l $HOME/tlog -s 38400 -c off -w 10 -t 150 < /dev/modem > /dev/modem &
# Open other clients here
suspend
!killall -KILL term

and finally, the script tstart which fires up the term clients is given by

#!/bin/sh
#
# This lets mail get out, can read news here, can pick up my mail here
#
/usr/local/bin/tredir 2025 25 2119 newshost:119 2110 pophost:110
#
# So I can open up Xwindows here
#
/usr/local/bin/trsh -s txconn
#
# So I will receive mail....
#

/usr/local/bin/pop
#
# Clean out the queue, in case of boo-boos
#
/usr/bin/runq
#
# Done now
#
echo ^G^G > /dev/console

When finally you want to close the connection, you resume and terminate kermit. The last line of the script kills the local term and returns the system in its initial state.

(Note of the author:  instead of doing '!killall -KILL term', I think it should be
possible just to do '!tshutdown'.  This should also work?)

As I said, there are zillions of ways to do this, these are just meant as examples to get you started.

8 Other programs and term

A number of programs can already be used in combination with term. To become these programs there are two possibilities:

1. Use ready made binaries.

2. Port and compile the sources.

8.1 Use ready made binaries

Some programs are modified to work with term. They include telnet, (nc)ftp, Mosaic and many others.

Binaries ofthese programs can be found on sunsite.unc.edu in /pub/Linux/apps/comm/termstuff and in
/pub/Linux/apps/comm/term/term/extra.

Most of these programs are compiled form term 1.17 or earlier. They should, however, still work with the newer versions of term as explained earlier.

9 Term speed table

Thanks to Bill McCarthy we now have a table with term speed information for different modems, term versions and connection conditions. Its purpose is to give new and experienced user some idea what other folks are using and the results they are getting.

                                LINUX TERM CHART 8/14/94

10. Term clients used in the world 22

|___modem speed/make___|___line speed__|__avg cps__|__high__|__term ver_|
| 1)  USR SP 14.4      |   9600        |  950      |  963   | 1.1.7     |
| 2)  USR SP 14.4      |  14400        | 1376      |  n/a   | 1.18p06   |
| 3)  Zoom 2400        |   2400        |  220      |  230   | 1.19      |
| 4)  Boca V.32bis 14  |  57600        | 1400      |  n/a   | 1.01/09?  |
| 5)  Viva 14.4        |  14400        | 1300      |  n/a   | 1.16      |
| 6)  USR SP 14.4      |  14400+       | 1550      | 1680   | 1.19      |
| 7)  Intel 14.4 Fax   |  14400        | 1400      | 1650   | 2.0.4     |
| 8)  cable tv hookup  |  57600        | 1500      | 1800   | 1.18p06   |
| 9)  Twincom 144/DFi  |  57600        | 1500      | 4000?  | 2.0.4     |
| 10) USR SP 14.4      |  14400        | 1200      | 1500   | 1.08      |
| 11) cable tv hookup  |  19200        | 1300      | 1800   | 1.1.9     |
|-----------------------------------------------------------------------|
+Command flags/termrc settings:
1) default escapes   2) window 5       3) baudrate 2400    4) n/a
   baudrate 9600        timeout 200       window 3
   window 10                              noise on
   timeout 150
5) compress off      6) baudrate 19200  7) ignore 19+17    8) compress off
   window 10            compress on        window 4           escape 0, 13,
   timeout 150                             timeout 90         16-19, 255
   baudrate 38400                                             baudrate 0
                                                              shift 224
                                                              flowcrtl 500
                                                              window 10
                                                              timeout 70
                                                              retrain on
                                                              breakout 24
9) compress off     10) compress off      11) baudrate 19200
   baudrate 57600       baudrate 38400        compress on
   window 10            escape 17, 19         shift 224
   timeout 200          remote                escape 0, 13 16-17
   noise on                                   19, 255
   share on                                   window 10
   remote                                     timeout 40
Escaping characters on one site also implies ignoring them on the other site.

10 Term clients used in the world

Below a list of application running with term is given. I am not stating that this list is complete so any completion is welcome. As far as possible I will present the site and directory where the application can be 11. Things to remember 23

found (all to my knowledge). When I state sunsite.unc.edu as place to find the application I mean that you can find it in one of the following two directories:

1. /pub/Linux/apps/comm/termstuff

2. /pub/Linux/apps/comm/term/term/extra

Here we go :-)

tupload             term package
trsh                term package
tmon                term package
tredir              term package
tudpredir           term package (versions 2.0.0 and higher)
txconn              term package
trdate(d)           term package
tshutdown           term package
libtermnet          term package
termftp
ncftp               sunsite.unc.edu
fsp                 sunsite.unc.edu
lynx                sunsite.unc.edu
Mosaic              sunsite.unc.edu:/pub/Linux/system/Network/Info-systems
chimera
finger              sunsite.unc.edu
rlogin
tin                 sunsite.unc.edu
slurp               sunsite.unc.edu
xgopher
gopher              sunsite.unc.edu
xboard              sunsite.unc.edu
ircII               sunsite.unc.edu
whois               sunsite.unc.edu
xwebster            sunsite.unc.edu
termtelnet          sunsite.unc.edu
term.mailerd+smail  sunsite.unc.edu
BCRMailHandlerXXX   bohr.physics.purdue.edu:/pub/bcr/term
sxpc                ftp.x.org:/R5contrib
xarchie             (compiled by some one, term version 2.0.x)
archie              (compiled by some one, term version 2.0.x)

11 Things to remember

In this section I try to present you a list of usefull ftp addresses, URL’s etc. where you can find software and information about term.

ftp:

sunsite.unc.edu:/pub/Linux/apps/comm/term/<whole-directory-tree>
sunsite.unc.edu:/pub/Linux/apps/comm/termstuff
sunsite.unc.edu:/pub/Linux/docs/HOWTO
bohr.physics.purdue.edu:/pub/bcr/term

URL:

http://sunsite.unc.edu/mdw/HOWTO/Term-HOWTO.html
http://zeus.cs.kun.nl:4080/term-howto/Term-HOWTO.html

netnews:

comp.os.linux.announce           announce of new Term-HOWTO
comp.os.linux.help               ask your questions about term here
comp.os.linux.misc               or here

When you start asking questions on netnews please make sure that you present people in the groups as much information as they need to solve your problem (term version number, way you set upyour connection etc.). At this moment many term versions are in use and all have their specific and common problems. Therefore, when you want a useful answer at least state the version of term you are using. Otherwise, in some cases only wild guesses are possible to help you save your problems.

12 Other Stuff

Some things that might be included:

• libtermnet.a

• Suggestions

Anyways, if you have comments, criticisms, suggestions, or anything else to say about this document, please fire away. As Bill Reynolds is very busy at this moment I, Patrick Reijnen, have taken over the the authorship of the Term-HOWTO. I can (currently) be reached at patrickr@sci.kun.nl or patrickr@cs.kun.nl.

13 Acks

Thanks go to a lot of people. First and foremost to Michael O’Reilly and all the developers of term, who have provided us with such a great tool. I would also like to thanks everyone who gave feedback and contributed to this HOWTO, they include Ronald Florence, Tom Payerle, Bill C. Riemers, Hugh Secker-Walker, Matt Welsh, Bill McCarthy, Sergio and everybody forgotten to mention.